solve the equation 2y^2-8y-5=0
you cant factor this out, i've tried to
im stuck on whats next, the home work says
1/2[2y^2-8y-5=0] then square both sides
why do you first use 1/2 with [2y^2-8y-5=0] ?

Answer by Alan3354(69443)

solve the equation 2y^2-8y-5=0
you cant factor this out, i've tried to
im stuck on whats next, the home work says
1/2[2y^2-8y-5=0] then square both sides
----------------
It might say multiply by 1/2, then complete the square of the eqn.
1/2[2y^2-8y-5=0]
--> y^2 - 4y = .25
Then add 4 (to both sides) to make the left side a square.
y^2 - 4y + 4 = 4.25
(y-2)^2 = 4.25
Now take sqrt of both sides
y-2 = sqrt(4.25)
y-2 = ~ 2.062
y = ~ 4.062
------------
y-2 = ~ -2.062
y = ~ 0.062